OpenStudy (anonymous):

(x^3y^2-y)dx+(x^2y^4-x)dy=0 solve using integrating factor

OpenStudy (anonymous):

Integrating factor is not apparent

OpenStudy (anonymous):

Please help

OpenStudy (paxpolaris):

no idea ... sry

OpenStudy (paxpolaris):

\[(x^3y^2-y)+(x^2y^4-x)dy/dx=0\]\[(x^3y^2-y)=\frac d {dx}\left(xy- \frac15x^2y^5\right)-y+\frac25xy^5\] ...no idea what i'm doing.... wolfram alpha doesn't show an elegant solution. http://www.wolframalpha.com/input/?i=%28x%5E3y%5E2-y%29dx%2B%28x%5E2y%5E4-x%29dy%3D0